Determine the given, formulate the null and alternative hypothesis in words and in

symbols, and the appropriate test statistic.


A seller claimed that her lip tint has a mean organic content of 90%. A rival seller asked 60 users of that lip tint and found that it has a mean organic content of 85% with a standard deviation of 5%. Test the claim at 1% level of significance and assume that the population is approximately normally distributed.


1
Expert's answer
2022-05-09T16:05:54-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=90"

"H_a:\mu\not=90"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=59" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.661759."

The rejection region for this two-tailed test is "R = \{t: |t| > 2.661759\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{85-90}{5/\sqrt{60}}\approx-7.746"


5. Since it is observed that "|t| =7.746>2.661759=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=59" degrees of freedom, "t=-7.746" is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 90, at the "\alpha = 0.01" significance level.